# Deblending the MIGHTEE-COSMOS survey with XID+: The resolved radio source counts to Jy

### v1

@@ 4.6.4 Goodness-of-Fit: The p-value Residual Statistic
To assess the goodness-of-fit for each source, we use the Bayesian p-value residual (Gelman1996) statistic, as defined in Shirley2021. This metric is a form of posterior predictive check that quantifies model fit to local data given the uncertainties. The resulting statistic is a probability, where a value of 0 indicates a model that always provides a good fit, and a value of 1 indicates a model that always provides a poor fit.
@@ 4.7 Recovery of Source Counts
A robust test of the entire process is to compare the Euclidean-normalized source counts () of the recovered catalog against the true input catalog to the simulation, where the injected source population is limited to flux densities above 0.5 Jy. Figure 14 shows the source counts. First, we examine the results from our PyBDSF blind source extraction. At bright flux densities ( mJy), the SRL counts are in excellent agreement with the true T-RECS counts. At intermediate flux densities ( mJy), the PyBDSF SRL catalog shows a significant upturn, a clear signature of flux boosting. Below mJy, the PyBDSF counts drop off sharply as the sources in this flux range are boosted to higher flux densities.
@@ 4.8 Summary and Recommended Strategy
This performance remains stable under realistic observational conditions; our stress tests with spatially varying noise and empirical positional offsets confirm that the strategy is robust to these real-world imperfections. Finally, the Bayesian p-value provides a necessary diagnostic for identifying the small fraction of poorly constrained sources. Based on these findings, our recommended strategy for the real MIGHTEE data is to use a high-confidence, Radio-Likely Prior list and apply a mask to exclude sources 50 Jy. We further recommend that the final catalog include the Bayesian p-value and detection significance to facilitate the selection of reliable samples.
@@ 5 Application to the MIGHTEE-COSMOS Field
Having established and tested our optimal deblending strategy on realistic simulations, including spatially varying noise and positional offsets, we now apply this approach to the MIGHTEE‑COSMOS low‑resolution map. This section details the construction of the final prior catalog and mask, and presents the results of the deblending; the outcome is a high-fidelity source catalog and number counts.
@@ 5.2.1 Radio Masking
Following the results from our simulations, our primary mask is generated from the PyBDSF SRL catalog by masking all sources brighter than 50 Jy. Specifically, we select all sources from the MIGHTEE DR1 SRL catalog from Hale2025 with a peak flux greater than 50 Jy and mask them.
Beyond the primary mask validated in simulations, we adopt a more aggressive elliptical mask around the very brightest sources ( mJy). This is necessary to address observational artifacts, such as residual sidelobes from imperfect deconvolution, which are not present in our idealized simulations. These artifacts can extend over larger areas than the source emission itself. We determined an empirical relationship between source flux and the required elliptical mask size by measuring the extent of these artifacts around several bright sources in the MIGHTEE-COSMOS map. This relationship defines progressively larger mask radii for sources in different flux density bins (Table 3). The final radio mask is the combination of the 50 Jy model-based mask and this additional artefact mask.
@@ 5.2.2 Stellar Masking
Finally, we mask regions around bright stars using the COSMOS2020 combined star mask (Weaver2022). We mask these regions because the multi-wavelength data used to define our prior are incomplete in these areas.
@@ 5.2.3 Final priors
Applying a combined mask, which incorporates both our radio masks and the star masks (Weaver2022), to our two raw prior lists yields the final input catalogs used for deblending. The Comprehensive LIR Prior is reduced to 298,894 sources. The more targeted Radio-Likely Prior is reduced to 89,562 sources.
@@ 5.3 Flux Recovery and Comparison
We ran XID+ on the MIGHTEE-COSMOS map using the Radio-Likely Prior, resulting in a full deblended catalog of 89,562 sources. The primary diagnostic for the reliability of individual flux measurements is the Bayesian p-value residual statistic. Our initial deblending run, using sampler settings identical to those validated in our simulations (i.e., the Stan default max_treedepth=10), led to a large number of sources with poor goodness-of-fit statistics (9,238 sources, or over 10 per cent of the sample, had a Bayesian p-value > 0.6).
@@
Further investigation showed that un-modeled artifacts and extended emission in the real MIGHTEE data produced complex posteriors that the sampler could not explore fully at the default depth, resulting in incomplete MCMC convergence. Increasing the max_treedepth parameter to 15 allowed the NUTS sampler to explore these posteriors more thoroughly. This computationally expensive change dramatically improved the fits, reducing the number of poorly fit sources by over 68 per cent (from 9,238 to 2,933), now representing just 3.3 per cent of the sample.
@@
To validate our flux densities, we compare them to the independent analysis of the MIGHTEE data using the super-deblended technique (Jin2018; Liu2018). This approach performs PSF-fitting at prior positions derived from deep ancillary data. We use the updated catalog (Jin2018; Jin2022; Jin2024), which was constructed using priors from the COSMOS2020 (Weaver2022), VLA 3GHz (Smolcic2017), and A3COSMOS (Liu2019) catalogs. This dataset serves as a robust benchmark because the super-deblended flux densities and errors were rigorously calibrated to quasi-Gaussian uncertainties using Monte Carlo simulations performed directly on the real map (Jin2018; Liu2018). This specific dataset has been used to derive radio properties in several recent studies (e.g., An2021; Delvecchio2021; Sillassen2024)
@@
We cross-matched our full Radio-Likely Prior catalog (89,562 sources) with the super-deblended catalog using a 0.3 arcsec radius, finding 88,939 matches. As shown in Figure 15, the two methods show excellent agreement for the bulk of the population, as they lie tightly along the 1:1 line with some scatter. The sources flagged with a high p-value in our analysis (dark green points) are mostly those that form the outlier population where our XID+ flux densities are systematically higher than the super-deblended values. This confirms that the p-value is an effective diagnostic for identifying potentially unreliable flux measurements.
@@
Our final data product is a complete catalog of 89,562 sources. For each source, we provide the Bayesian p-value statistic, which assesses the goodness of fit, and the detection significance. We do not remove sources from the catalog, as the full posterior distributions for low-significance sources remain statistically valid for stacking or ensemble analyses.
@@
However, for applications requiring robust individual source detections, we first consider a simple significance‑based selection. Applying the criterion defined in Section 4.6.3 (Significance ) yields 31,779 sources, while a flux density cut above the average limit (Jy) yields 31,083 sources. To construct a highly conservative recommended sample, we select sources that satisfy both of these criteria simultaneously and exhibit a reliable goodness-of-fit (p-value residual statistic ). This rigorous selection results in a high-fidelity catalog of 20,757 sources.
@@ 5.5 Radio Source Number Counts
The primary scientific product of this work is the 1.4 GHz radio source number counts. The flux densities have been converted from the native MIGHTEE observing frequency to 1.4 GHz using the spatially varying frequency-response maps of Hale2025, adopting a spectral index . In Figure 17, we present the Euclidean-normalized source counts derived from our Comprehensive and Radio-Likely Prior catalogs. Our primary result, from the Radio-Likely Prior, shows excellent agreement with the literature at certain flux ranges. At the bright end of our deblended range, from approximately 20 to 40 Jy, our counts lie below the completeness-corrected literature values from Hale2025. This turnover is expected from the 50 Jy mask and a similar behavior is seen in the simulations (see Figure 14), confirming that this is a feature of our methodology, not a discrepancy in the measurement.
Moving deeper into the confusion-dominated regime in MIGHTEE, from approximately 15 Jy down to our faintest reliable bin at 4.8 Jy (3), our counts show excellent agreement with the P(D) analysis of the MeerKAT DEEP2 field from Matthews2021. The agreement between our catalog-based deblended counts and this independent statistical measurement validates our methodology in a flux density regime where traditional source-finding fails due to confusion. This implies that we are resolving the vast majority of the radio background at these flux densities.
Below this 3 limit, our counts decline and fall below the P(D) result. This decline is expected since, based on the analysis of the simulations, below 3 the flux densities of the Radio-Likely Prior source counts systematically and increasingly underestimate the ground truth. Although we note that the only observational constraints on the true form of the source counts at these flux densities come from the P(D) analysis of Matthews2021, where optical counterparts are not required.
@@ 6 Conclusions and future work
Deep radio continuum surveys from highly sensitive facilities like MeerKAT provide unprecedented access to the faint radio source population, enabling important tests of galaxy evolution models. However, their scientific potential is fundamentally limited by source confusion in the absence of appropriate analysis techniques. We have developed, validated, and applied a complete framework for producing high-fidelity deblended radio source catalogs from confusion-limited survey data, using the MIGHTEE-COSMOS field as our primary application. Our key findings and contributions are:
@@
This work delivers the deepest catalog-based MIGHTEE-COSMOS source counts to date. When applied to the MIGHTEE-COSMOS field, our validated strategy recovers the 1.4 GHz source number counts down to the limit (Jy). The final catalog contains 89,562 sources, including a recommended high-fidelity subset of 20,757 sources.
@@
Bayesian diagnostics and sampler optimization are important for ensuring catalog fidelity: The Bayesian p-value residual statistic is a powerful tool for identifying sources with potentially unreliable flux measurements. We find that its successful application to complex, real-world data requires careful tuning of the MCMC sampler settings within the Stan framework. Increasing the max_treedepth from the default of 10 to 15 was required for convergence on robust solutions, though with increased computational cost. This demonstrates that the increased complexity of real data requires a more exhaustive exploration of the posterior parameter space to achieve reliable fits.
@@
This deblending framework enables a wide range of science applications beyond what is possible with traditional source-finding. These include constructing a complete, dust-unbiased census of the cosmic star formation history (e.g. Novak2017; Matthews2024); probing the evolution of the infrared–radio correlation to unprecedented depths (e.g. Delhaize2017; Delvecchio2021; Moloko2025); disentangling star formation and AGN activity in radio-quiet AGN (e.g. Padovani2016; Panessa2019) and precisely measuring the faint-end slope of the radio luminosity function (e.g. Smolcic2017; Algera2020). The framework established in this work is directly applicable to other confusion-limited radio surveys from current facilities (ASKAP, LOFAR, VLA, uGMRT).
@@
While the framework established in this work provides a validated approach for producing deep, reliable deblended radio source catalogs, it could be refined further.
@@
Specifically, in the current framework, we assume that the sources that lie at these faint flux densities are both unresolved (a reasonable assumption) and that their flux density is distributed over the restoring beam. However, the maps that we use are not cleaned to the levels that we are extracting the flux densities from, as such, we should use the synthesised beam as our PSF rather than the restoring beam, whose information is in the headers. However, this would introduce additional complexity without a significant amount of gain. The synthesised beam for the MIGHTEE data is tapered to produce a Gaussian-like profile that is very similar to the adopted restoring beam. As such, the difference between using the restoring beam approximation and the full synthesised beam is of the order 5 per cent, significantly less than the uncertainties on individual sources. However, ideally, one would simulate the full observational pipeline: Using the improved sky model, one could simulate the entire data acquisition and reduction process. This involves generating the raw instrumental visibilities, introducing realistic thermal noise and calibration errors, and then processing these data through the identical imaging and deconvolution pipeline used for the real MIGHTEE observations. This approach would naturally produce complex artifacts, such as residual sidelobes and spatially correlated noise, providing the ultimate test of this methodology against real-world instrumental effects. However, this would be very computationally intensive.
@@
One could also improve the Radio-Likely Prior with more advanced source selection. The success of our deblending is highly dependent on the purity of the Radio-Likely Prior. While effective, the current method based on the IRRC can be advanced by using more sophisticated predictive models. Future work will explore the use of full-SED models to measure the SFR and/or machine learning algorithms trained on the rich, multi-wavelength photometry available in the COSMOS field, similar to (Wang2024b) to learn the complex relationship between a galaxy’s full UV-to-mid-infrared spectral energy distribution and its expected 1.4 GHz flux. This would allow us to generate a more nuanced, probabilistic radio flux estimate for every source in the parent catalog, creating a more robust and complete Radio-Likely Prior that better distinguishes detectable radio sources from undetectable ones.
@@ Acknowledgements
EM acknowledges financial support from the University of South Africa (Research Fund: 409000). MJJ, CLH and IHW acknowledges support from the Hintze Family Charitable Foundation through the Oxford Hintze Centre for Astrophysical Surveys. MJJ acknowledges the support of the STFC consolidated grant [ST/S000488/1] and [ST/W000903/1] and from a UKRI Frontiers Research Grant [EP/X026639/1]. MGS acknowledges support from the South African National Research Foundation (Grant No. 84156). DJBS acknowledges support from the UK’s Science and Technology Facilities Council via grant ST/Y001028/1, and from the Leverhulme Trust via Research Project Grant RPG-2025-078. CLH acknowledges support from STFC through grant ST/Y000951/1. MV acknowledges financial support from the Inter-University Institute for Data Intensive Astronomy (IDIA), a partnership of the University of Cape Town, the University of Pretoria and the University of the Western Cape, and from the South African Department of Science and Innovation’s National Research Foundation under the ISARP RADIOMAP Joint Research Scheme (DSI-NRF Grant Number 150551) and the CPRR HIPPO Project (DSI-NRF Grant Number SRUG22031677). SJ acknowledges the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie Action grant No. 101060888, and the Villum Fonden research grants 37440 and 13160. FXA acknowledges the support from the National Natural Science Foundation of China (12303016) and the Natural Science Foundation of Jiangsu Province (BK20242115). The authors thank Suman Chatterjee, Antonio la Marca, and Lingyu Wang for helpful discussions that enhanced this work.
@@
The MIGHTEE DR1 data products are available at https://doi.org/10.48479/7msw-r692. The final deblended catalog presented in this paper will be made available upon publication.

### v2

@@ 4.6.4 Goodness-of-Fit: The p-value Residual Statistic
To assess the goodness-of-fit for each source, we use the Bayesian p-value residual (Gelman et al., 1996) statistic, as defined in Shirley et al. (2021). This metric is a form of posterior predictive check that quantifies model fit to local data given the uncertainties. The resulting statistic is a probability, where a value of 0 indicates a model that always provides a good fit, and a value of 1 indicates a model that always provides a poor fit.
@@ 4.7 Recovery of Source Counts
A robust test of the entire process is to compare the Euclidean-normalized source counts () of the recovered catalogue against the true input catalogue to the simulation, where the injected source population is limited to flux densities above 0.5 Jy. Figure 14 shows the source counts. First, we examine the results from our PyBDSF blind source extraction. At bright flux densities ( mJy), the SRL counts are in excellent agreement with the true T-RECS counts.
At intermediate flux densities ( mJy), the PyBDSF SRL catalogue shows a significant upturn, a clear signature of flux boosting. Below mJy, the PyBDSF counts drop off sharply as the sources in this flux range are boosted to higher flux densities.
@@ 4.8 Summary and Recommended Strategy
This performance remains stable under realistic observational conditions; our stress tests with spatially varying noise and empirical positional offsets confirm that the strategy is robust to these real-world imperfections. Finally, the Bayesian p-value provides a necessary diagnostic for identifying the small fraction of poorly constrained sources. Based on these findings, our recommended strategy for the real MIGHTEE data is to use a high-confidence, Radio-Likely Prior list and apply a mask to exclude sources 50 Jy. We further recommend that the final catalogue include the Bayesian p-value and detection significance to facilitate the selection of reliable samples.
@@ 5 Application to the MIGHTEE-COSMOS Field
Having established and tested our optimal deblending strategy on realistic simulations, including spatially varying noise and positional offsets, we now apply this approach to the MIGHTEE‑COSMOS low‑resolution map. This section details the construction of the final prior catalogue and mask, and presents the results of the deblending; the outcome is a high-fidelity source catalogue and number counts.
@@ 5.2.1 Radio Masking
Following the results from our simulations, our primary mask is generated from the PyBDSF SRL catalogue by masking all sources brighter than 50 Jy. Specifically, we select all sources from the MIGHTEE DR1 SRL catalogue from Hale et al. (2025) with a peak flux greater than 50 Jy and mask them.
Beyond the primary mask validated in simulations, we adopt a more aggressive elliptical mask around the very brightest sources ( mJy). This is necessary to address observational artefacts, such as residual sidelobes from imperfect deconvolution, which are not present in our idealized simulations. These artefacts can extend over larger areas than the source emission itself. We determined an empirical relationship between source flux and the required elliptical mask size by measuring the extent of these artefacts around several bright sources in the MIGHTEE-COSMOS map. This relationship defines progressively larger mask radii for sources in different flux density bins (Table 3). The final radio mask is the combination of the 50 Jy model-based mask and this additional artefact mask.
@@ 5.2.2 Stellar Masking
Finally, we mask regions around bright stars using the COSMOS2020 combined star mask (Weaver et al., 2022). We mask these regions because the multi-wavelength data used to define our prior are incomplete in these areas.
@@ 5.2.3 Final priors
Applying a combined mask, which incorporates both our radio masks and the star masks (Weaver et al., 2022), to our two raw prior lists yields the final input catalogues used for deblending. The Comprehensive LIR Prior is reduced to 298,894 sources. The more targeted Radio-Likely Prior is reduced to 89,562 sources.
@@ 5.3 Flux Recovery and Comparison
We ran XID+ on the MIGHTEE-COSMOS map using the Radio-Likely Prior, resulting in a full deblended catalogue of 89,562 sources. The primary diagnostic for the reliability of individual flux measurements is the Bayesian p-value residual statistic. Our initial deblending run, using sampler settings identical to those validated in our simulations (i.e., the Stan default max_treedepth=10), led to a large number of sources with poor goodness-of-fit statistics (9,238 sources, or over 10 per cent of the sample, had a Bayesian p-value > 0.6).
@@
Further investigation showed that un-modeled artefacts and extended emission in the real MIGHTEE data produced complex posteriors that the sampler could not explore fully at the default depth, resulting in incomplete MCMC convergence. Increasing the max_treedepth parameter to 15 allowed the NUTS sampler to explore these posteriors more thoroughly. This computationally expensive change dramatically improved the fits, reducing the number of poorly fit sources by over 68 per cent (from 9,238 to 2,933), now representing just 3.3 per cent of the sample.
@@
To validate our flux densities, we compare them to the independent analysis of the MIGHTEE data using the super-deblended technique (Jin et al., 2018; Liu et al., 2018). This approach performs PSF-fitting at the (super-deblended) prior positions described in Section 2.4.3. The resulting catalogue (Sillassen et al., 2024) provides flux densities and errors that were rigorously calibrated to quasi-Gaussian uncertainties using Monte Carlo simulations performed directly on the real map (Jin et al., 2018; Liu et al., 2018). This catalogue has been widely used in many works (e.g., An et al., 2021; Delvecchio et al., 2021; Jin et al., 2022, 2024; Sillassen et al., 2024).
@@
We cross-matched our full Radio-Likely Prior catalogue (89,562 sources) with the super-deblended catalogue using a 0.3 arcsec radius, finding 88,939 matches. As shown in Figure 15, the two methods show excellent agreement for the bulk of the population, as they lie tightly along the 1:1 line with some scatter. The sources flagged with a high p-value in our analysis (dark green points) are mostly those that form the outlier population where our XID+ flux densities are systematically higher than the super-deblended values. This confirms that the p-value is an effective diagnostic for identifying potentially unreliable flux measurements.
@@
Our final data product is a complete catalogue of 89,562 sources. For each source, we provide the Bayesian p-value statistic, which assesses the goodness of fit, and the detection significance. We do not remove sources from the catalogue, as the full posterior distributions for low-significance sources remain statistically valid for stacking or ensemble analyses.
@@
However, for applications requiring robust individual source detections, we first consider a simple significance‑based selection. Applying the criterion defined in Section 4.6.3 (Significance ) yields 31,779 sources, while a flux density cut above the average limit (Jy) yields 31,083 sources. To construct a highly conservative recommended sample, we select sources that satisfy both of these criteria simultaneously and exhibit a reliable goodness-of-fit (p-value residual statistic ). This rigorous selection results in a high-fidelity catalogue of 20,757 sources.
@@ 5.5 Radio Source Number Counts
The primary scientific product of this work is the 1.4 GHz radio source number counts. The flux densities have been converted from the native MIGHTEE observing frequency to 1.4 GHz using the spatially varying frequency-response maps of Hale et al. (2025), adopting a spectral index . In Figure 17, we present the Euclidean-normalized source counts derived from our Comprehensive and Radio-Likely Prior catalogues. Our primary result, from the Radio-Likely Prior, shows excellent agreement with the literature at certain flux ranges. At the bright end of our deblended range, from approximately 20 to 40 Jy, our counts lie below the completeness-corrected literature values from Hale et al. (2025). This turnover is expected from the 50 Jy mask and a similar behavior is seen in the simulations (see Figure 14), confirming that this is a feature of our methodology, not a discrepancy in the measurement.
Moving deeper into the confusion-dominated regime in MIGHTEE, from approximately 15 Jy down to our faintest reliable bin at 4.8 Jy (3), our counts show excellent agreement with the P(D) analysis of the MeerKAT DEEP2 field from Matthews et al. (2021b). The agreement between our catalog-based deblended counts and this independent statistical measurement validates our methodology in a flux density regime where traditional source-finding fails due to confusion. This implies that we are resolving the vast majority of the radio background at these flux densities.
Below this 3 limit, our counts decline and fall below the P(D) result. This decline is expected since, based on the analysis of the simulations, below 3 the flux densities of the Radio-Likely Prior source counts systematically and increasingly underestimate the ground truth. Although we note that the only observational constraints on the true form of the source counts at these flux densities come from the P(D) analysis of Matthews et al. (2021b), where optical counterparts are not required.
@@ 6 Conclusions and Future Work
Deep radio continuum surveys from highly sensitive facilities like MeerKAT provide unprecedented access to the faint radio source population, enabling important tests of galaxy evolution models. However, their scientific potential is fundamentally limited by source confusion in the absence of appropriate analysis techniques. We have developed, validated, and applied a complete framework for producing high-fidelity deblended radio source catalogues from confusion-limited survey data, using the MIGHTEE-COSMOS field as our primary application. Our key findings and contributions are:
@@
This work delivers the deepest catalog-based MIGHTEE-COSMOS source counts to date. When applied to the MIGHTEE-COSMOS field, our validated strategy recovers the 1.4 GHz source number counts down to the limit (Jy). The final catalogue contains 89,562 sources, including a recommended high-fidelity subset of 20,757 sources.
@@
Bayesian diagnostics and sampler optimization are important for ensuring catalogue fidelity: The Bayesian p-value residual statistic is a powerful tool for identifying sources with potentially unreliable flux measurements. We find that its successful application to complex, real-world data requires careful tuning of the MCMC sampler settings within the Stan framework. Increasing the max_treedepth from the default of 10 to 15 was required for convergence on robust solutions, though with increased computational cost. This demonstrates that the increased complexity of real data requires a more exhaustive exploration of the posterior parameter space to achieve reliable fits.
@@
This deblending framework enables a wide range of science applications beyond what is possible with traditional source-finding. These include constructing a complete, dust-unbiased census of the cosmic star formation history (e.g.,Novak et al. 2017; Matthews et al. 2024); probing the evolution of the infrared–radio correlation to unprecedented depths (e.g.,Delhaize et al. 2017; Delvecchio et al. 2021; Moloko et al. 2025); disentangling star formation and AGN activity in radio-quiet AGN (e.g.,Padovani 2016; Panessa et al. 2019) and precisely measuring the faint-end slope of the radio luminosity function (e.g.,Smolčić et al. 2017; Algera et al. 2020). The framework established in this work is directly applicable to other confusion-limited radio surveys from current facilities (ASKAP, LOFAR, VLA, uGMRT).
@@
While the framework established in this work provides a validated approach for producing deep, reliable deblended radio source catalogues, it could be refined further.
@@
Specifically, in the current framework, we assume that the sources that lie at these faint flux densities are both unresolved (a reasonable assumption) and that their flux density is distributed over the restoring beam. However, the maps that we use are not cleaned to the levels that we are extracting the flux densities from, as such, we should use the synthesised beam as our PSF rather than the restoring beam, whose information is in the headers. However, this would introduce additional complexity without a significant amount of gain. The synthesised beam for the MIGHTEE data is tapered to produce a Gaussian-like profile that is very similar to the adopted restoring beam. As such, the difference between using the restoring beam approximation and the full synthesised beam is of the order 5 per cent, significantly less than the uncertainties on individual sources. However, ideally, one would simulate the full observational pipeline: Using the improved sky model, one could simulate the entire data acquisition and reduction process. This involves generating the raw instrumental visibilities, introducing realistic thermal noise and calibration errors, and then processing these data through the identical imaging and deconvolution pipeline used for the real MIGHTEE observations. This approach would naturally produce complex artefacts, such as residual sidelobes and spatially correlated noise, providing the ultimate test of this methodology against real-world instrumental effects. However, this would be very computationally intensive.
@@
One could also improve the Radio-Likely Prior with more advanced source selection. The success of our deblending is highly dependent on the purity of the Radio-Likely Prior. While effective, the current method based on the IRRC can be advanced by using more sophisticated predictive models. Future work will explore the use of full-SED models to measure the SFR and/or machine learning algorithms trained on the rich, multi-wavelength photometry available in the COSMOS field, similar to (Wang et al., 2024b) to learn the complex relationship between a galaxy’s full UV-to-mid-infrared spectral energy distribution and its expected 1.4 GHz flux. This would allow us to generate a more nuanced, probabilistic radio flux estimate for every source in the parent catalogue, creating a more robust and complete Radio-Likely Prior that better distinguishes detectable radio sources from undetectable ones.
@@ Acknowledgements
We thank the referee for useful suggestions which improved the quality of the paper. EM acknowledges financial support from the University of South Africa (Research Fund: 409000). MJJ, CLH and IHW acknowledges support from the Hintze Family Charitable Foundation through the Oxford Hintze Centre for Astrophysical Surveys. MJJ acknowledges the support of the STFC consolidated grant [ST/S000488/1] and [ST/W000903/1] and from a UKRI Frontiers Research Grant [EP/X026639/1]. MGS acknowledges support from the South African National Research Foundation (Grant No. 84156). DJBS acknowledges support from the UK’s Science and Technology Facilities Council via grant ST/Y001028/1, and from the Leverhulme Trust via Research Project Grant RPG-2025-078. CLH acknowledges support from STFC through grant ST/Y000951/1. JA acknowledges financial support from the Science and Technology Foundation (FCT, Portugal) through research grants UIDB/04434/2020 (DOI: 10.54499/UIDB/04434/2020), UIDP/04434/2020 (DOI: 10.54499/UIDP/04434/2020) and UID/04434/2025. MV acknowledges financial support from the Inter-University Institute for Data Intensive Astronomy (IDIA), a partnership of the University of Cape Town, the University of Pretoria and the University of the Western Cape, and from the South African Department of Science and Innovation’s National Research Foundation under the ISARP RADIOMAP Joint Research Scheme (DSI-NRF Grant Number 150551) and the CPRR HIPPO Project (DSI-NRF Grant Number SRUG22031677). SJ acknowledges the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie Action grant No. 101060888, and the Villum Fonden research grants 37440 and 13160. FXA acknowledges the support from the National Natural Science Foundation of China (12303016) and the Natural Science Foundation of Jiangsu Province (BK20242115). The authors thank Suman Chatterjee, Antonio la Marca, and Lingyu Wang for helpful discussions that enhanced this work.
@@
The MIGHTEE DR1 data products are available at https://doi.org/10.48479/7msw-r692. The deblended radio source catalogues (the Radio-Likely and Comprehensive samples) generated in this work are available in a VizieR catalogue available at CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/MNRAS.